A car drives straight off the edge of a cliff that is 54 m high. The police at the scene of the accident note
that the point of impact is 130 m from the base of the cliff. How fast was the car traveling when it went
over the cliff?
1
Expert's answer
2020-11-10T06:56:58-0500

Let's write the equations of motion of the car in horizontal and vertical directions:


"x=v_0t, (1)""y=\dfrac{1}{2}gt^2, (2)"

here, "x=130\ m" is the horizontal displacement of the car (or the distance from the base of the cliff), "v_0" is the initial velocity of the car when it went over the cliff, "t" is the time that the car takes to reach the ground, "y=54\ m" is the vertical displacement of the car (or the height of the cliff), "g = 9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the car takes to reach the ground from the second equation:


"t = \sqrt{\dfrac{2y}{g}}."

Then, we can substitute "t" into the first equation and find "v_0":


"x=v_0\sqrt{\dfrac{2y}{g}},""v_0=\dfrac{x}{\sqrt{\dfrac{2y}{g}}}=\dfrac{130\ m}{\sqrt{\dfrac{2\cdot 54\ m}{9.8\ \dfrac{m}{s^2}}}}=39\ \dfrac{m}{s}."

Answer:

"v_0=39\ \dfrac{m}{s}."